1.  A firm raises capital by selling $20,000 worth of debt with flotation costs equal to 2% of its par value. If the debt matures in 10 years and has a coupon interest rate of 8%, what is the bond's YTM?


1
Expert's answer
2021-11-14T17:36:50-0500

Par value of the bond = $20,000

Present value of bond = $20,000 * 2% = $19,600

Number of years to maturity of the bonds = 10 years

Coupon rate = 8%

Coupon amount = 8%*$20,000 = $1,600

"Bond \; YTM = \frac{Coupon + \frac{Face \;Value - Price }{Number \;of \;years \;to \;maturity}}{\frac{Face \;Value + Price}{2}}\\

=\frac{1600 + \frac{20000 - 19600 }{10}}{\frac{19600 + 20000}{2}}\\

=\frac{1600+40}{19800}\\

=0.0828282"

YTM=8.28%